OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (anonymous):

It's 0 hun.

OpenStudy (anonymous):

x = 0.

OpenStudy (uri):

8x=-8+5+3 8x=16 x=16/8 x=2 ans

OpenStudy (souvik):

-8+5+3=?

OpenStudy (uri):

0/8? I'm confused

OpenStudy (goformit100):

@uri which is correct -2 or 0 ??

OpenStudy (uri):

0.

OpenStudy (anonymous):

0 is correct!

OpenStudy (uri):

0/8 =0 right? :)

OpenStudy (anonymous):

It's an easy problem they cancel out. Therefore it's 0.

OpenStudy (uri):

Yea @mercedesjaay_ but you directly gave the answer which ain't allowed here. )

OpenStudy (anonymous):

Well, you needed the answer to your problem therefore I gave you the answer, so you can't complain there. Hope I helped, goodbye.

OpenStudy (uri):

My problem? :3

OpenStudy (anonymous):

@uri You directly gave her the answer too! When you said "0." So wth?

OpenStudy (uri):

Oh if you don't see i went step by step maybe you need to scroll up? You just said *it's 0 hun*

OpenStudy (anonymous):

Yeah, and below you gave her the wrong answer.. Probably on purpose to but whatever. @ysenia it's 0. :)

OpenStudy (uri):

Nvm